Circular No. 6088

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994aa IN NGC 1320 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova of mag 17 at α = 3^{h}22^{m}18^{s}.48, δ = −3^{o}13′05′′.7 (equinox 1950.0), 10′′.1 east and 1′′.8 north of the center of NGC 1320. The object was detected on a U.K. Schmidt I plate taken by C. P. Cass on Sept. 11 and confirmed on a CCD image taken by McNaught and G. J. Garradd with the Siding Spring Observatory’s 1.0-m reflector on Sept. 17. No image appears at the position on a J survey plate taken in 1982.

KS 1730−312

K. Borozdin, N. Alexandrovich and R. Sunyaev, Space Research Institute, Moscow, on behalf of the MIR-KVANT team, report: “This x-ray transient (cf. IAUC 6083) continues to increase in luminosity in the standard x-ray spectral band. According to data obtained on Sept. 26 with the MIR-KVANT-TTM coded-mask instrument, the spectrum can be approximated by a power law with photon index 2.7 ± 0.2 and column density N_{H} = 4 × 10^{22} cm^{−2}. The flux at 4 keV was ∼ 600 mCrab. An even higher flux (850 mCrab at 4 keV) and softer spectrum were detected on Sept. 28, when the power-law index was 4.2 ± 0.2 and N_{H} = 8 × 10^{22} cm^{−2}. Possibly this is another x-ray nova.”

JUPITER AND COMET 1993e O. P. N. Calla, V. Sukumaran, B. M. Darji, M. R. Deshpande, H. O.

Vats, P. Janardan and N. M. Vadher, Space Applications Center and Physical Research Laboratory, Ahmedabad, report: “Observations of Jupiter before, during and after the impacts of fragments of comet 1993e made at 4.15 GHz using a 14-m dish at the Ahmedabad Earth Station indicate that the microwave radio emission steadily increased after the impact of fragment A on July 16. This increase in continuum emission was ∼ 100 percent, and on July 25 the flux returned to pre-impact level. In addition, the four impacts K, N, P_{2} and S showed large bursts of microwave emission of the order of 5.1, 5.2, 4.3 and 2.3 times, respectively. These bursts were observed with a time delay of ∼ 30 min. The duration of these bursts ranged from 19 to 36 min. The digitally recorded data have a time resolution of 48 ms and are being analyzed in detail.”

1994 October 4 Brian G. Marsden